### v1

# Studying the Impact of Early Test Termination Due to Assertion Failure on Code Coverage and Spectrum-based Fault Localization

Abstract
An assertion is commonly used to validate the expected program’s behavior (e.g., if the returned value of a method equals an expected value) in software testing. Although it is a recommended practice to use only one assertion in a single test to avoid code smells (e.g., Assertion Roulette), it is common to have multiple assertions in a single test. One issue with tests that have multiple assertions is that when the test fails at an early assertion (not the last one), the test will terminate at that point, and the remaining testing code will not be executed. This, in turn, can potentially reduce the code coverage and the performance of techniques that rely on code coverage information (e.g., spectrum-based fault localization). We refer to such a scenario as early test termination. Understanding the impact of early test termination on test coverage is important for software testing and debugging, particularly for the techniques that rely on coverage information obtained from the testing. We conducted the first empirical study on early test termination due to assertion failure (i.e., early test termination) by investigating 207 versions of 6 open-source projects. We found that a non-negligible portion of the failed tests (19.1%) is early terminated due to assertion failure. Our findings indicate that early test termination harms both code coverage and the effectiveness of spectrum-based fault localization. For instance, after eliminating early test termination, the line/branch coverage is improved in 55% of the studied versions, and improves the performance of two popular SBFL techniques Ochiai and Tarantula by 15.1% and 10.7% compared to the original setting (without eliminating early test termination) in terms of MFR, respectively.
Index Terms:
Early Test Termination, Coverage, Fault Localization, Assertion FailureI Introduction
Testing is an effective and practical way of finding bugs and improving software quality. However, due to the scale of modern software and the complexity of tests, it can be difficult for developers to know the cause of test failures and propose fixes. Prior studies have proposed various automated testing techniques that leverage code coverage and test execution information to assist developers with test failure diagnosis or repair. For example, researchers have proposed using code coverage information in failed/passed tests for tasks such as spectrum-based fault localization (SBFL) [1, 2, 3, 4, 5, 6, 7, 8] and automated program repair [9, 10, 11, 12, 13]. For instance, Ochiai [14] and Tarantula [15], SBFL techniques, leverage both passed/failed tests together with their coverage information to identify the faulty location in the program. In modern continuous integration pipelines, SBFL is deployed and triggered automatically when a test fails [16, 17]. Therefore, when assertion failure happens, SBFL provides additional information that helps developers narrow the search space and identify the faulty code effectively.
In testing, assertions are commonly used to verify the program’s expected behavior (e.g., if the returned value of a method equals an expected value). Although it is a recommended practice to use only one assertion in a test to avoid code smell (e.g., Assertion Roulette [18]), prior studies [19, 20] found that many tests may still contain multiple assertions. One issue with tests that have more than one assertion is that when the test fails at an early assertion, the test will terminate at that point and skip the remaining test code. This, in turn, can potentially reduce the code coverage, hindering the effectiveness of SBFL or automated program repair techniques. We refer to the scenario where a test has multiple assertions and fails at an early assertion (i.e., not the last assertion) as early test termination. Understanding the impact of early test termination on test coverage is important for software testing and debugging, particularly for the techniques that rely on coverage information obtained from the testing.
Therefore, we first conduct an empirical study to understand the prevalence of early test termination and examine its impact on the test coverage and effectiveness of SBFL techniques (i.e., Ochiai and Tarantula) that rely on coverage information. We investigated 207 versions collected from six open-source projects used in Defect4J [21] and T-Evos [1]. We formulate the following research questions in our research:
RQ1: How prevalent is early test termination?
We investigate how prevalent early test termination is in failed tests and the amount of test code that is skipped due to early test termination.Results: Among the failed tests, a non-negligible portion (19.1%) is early terminated due to assertion failure, and they occur in 29% of studied versions. Due to early test termination, on average 15.3% to 60.5% of the test code is skipped.
RQ2: How does early test termination affect test coverage?
We apply two approaches Trycatch and Slicing to enforce the continuous execution of tests even if early test termination happens. In Trycatch, we modify a test by adding a try-catch block surrounding each assertion in the test to enforce it to keep executing. In Slicing, we use program slicing to disassemble a test with multiple assertions into multiple independent sub-tests, which contain only one assertion and its dependent code. We refer to the original setting as Original. We compare the testing coverage before (Original) and after applying Trycatch and Slicing.Results: In general, early test termination negatively impacts the coverage in the subject under test in terms of all granularities (i.e., instruction, line, branch, and method). Trycatch and Slicing improve the line/branch coverage in 55% of the studied versions compared with the Original.
RQ3: How does early test termination affect the performance of spectrum-based fault localization?
To evaluate the impact of early test termination on SBFL, we select two popular techniques Ochiai and Tarantula as the representatives of SBFL techniques. We compare the performance of Ochiai and Tarantula before (Original) and after applying Trycatch and Slicing, to investigate the impact of early test termination on SBFL.Results: early test termination negatively impacts the effectiveness of SBFL. Both Slicing and Trycatch improve the performance of SBFL by eliminating early test termination. Slicing outperforms Trycatch for SBFL. On average, Slicing improves the performance of Ochiai and Tarantula by 15.1% and 10.7% compared with Original in terms of Mean First Rank (MFR), respectively.
We summarize our contribution as follows:
We conducted the first empirical study on early test termination due to assertion failure (i.e., early test termination). We provide evidence of the prevalence of early test termination and its negative impact on testing coverage and the effectiveness of SBFL.
We evaluate using two approaches (Trycatch and Slicing) to mitigate the impact of early test failures on test coverage and demonstrate that Slicing performs better.
We provide actionable suggestions to prevent early test termination, and approaches to mitigate early test termination if it already exists in their project.
We make our dataset publicly available [22] to encourage future research on this direction.
II Background
In this section, we first discuss the phenomenon of early test termination and its potential impact on SBFL. Then, we provide a motivating example to illustrate the impact of early test termination.
II-A Early Test Termination
To assist developers with diagnosing the causes of test failures, researchers have proposed various techniques such as SBFL [1, 2, 3, 4, 6, 7, 8]. However, when a test terminates prematurely, the generated test coverage and test failure messages become incomplete, which affects how developers and techniques such as SBFL locate the root cause of failures.
We first formally define early test termination. Suppose there is a test that has statements {, , …, ,…, }. We define as an early terminated test if terminates at a statement , where . A test may terminate early due to various reasons, such as errors in object initialization, method invocation, or assertion. In this paper, we focus the study on early test termination due to assertion failure. Although having multiple assertions in a test can be considered a test smell [18], we observed that this is a common practice in the studied systems (as shown in Table I, 34.6% to 73.3% of the tests have multiple assertions in our studied projects). We also observe that early test termination due to assertion failure happens frequently (see more details in Section IV). In JUnit, if an assertion fails, the test will terminate immediately at that point, skipping the code that follows, and it will report the test as failed. For instance, in Figure LABEL:lst:example1, the test testRootEndpoints has multiple assertions, and it stops at the third assertion. As a result, the remaining test code after the failed assertion would not be executed, which potentially leads to incomplete coverage of the subject under test. Our goal is to study the impact of early test termination due to assertion failure on test coverage and the performance of SBFL. For simplicity, we refer to early test failure due to assertion failure as early test termination and the test where early test termination happens as Testearly in the rest of the paper.
II-B Spectrum-based Fault Localization
In this work, we focus on Spectrum-Based Fault Localization (SBFL) as the subject technique to understand the impact of early test termination. We choose SBFL because it relies directly on the quality and completeness of the collected code coverage, and it is an essential step for automated program repair. SBFL uses statistical formulas to measure the suspiciousness of program units based on the program execution traces. Execution traces, also called program spectra, contain details of the covered statements in failed and passed tests. The suspiciousness scores computed from program spectra are then used to generate a ranked list of program units that are most likely to be responsible for the fault. A number of SBFL techniques have been developed over the decades, such as Tarantula [15], Ochiai [14], DStar [8], and BARINEL [6]. In particular, Ochiai and Tarantula are commonly used SBFL techniques because of their high effectiveness [2, 23]. For instance, Ochiai is calculated as: . The suspiciousness score of a statement is based on the execution of passed and failed tests. Therefore, the performance of SBFL is potentially impacted by early test termination. For instance, suppose a test has two assertions and to test two branches and in the subject under test, respectively. If the test fails at , then only is covered. If no other tests cover , the Ochiai score for all the statements in will be 0. However, may also fail if the test were able to continue executing, and we would miss the fault in because of early test termination.
III Experimental Design
In this section, we first present our research questions (RQs). We then present the data collection process and our approach to answering the research questions.
III-A Research Questions
The goal of our study is twofold: 1) understand the prevalence of early test termination; 2) study the impact of early test termination on the test coverage and performance of debugging techniques that rely on test coverage (SBFL in this study). Therefore, we formulate our research to answer the following research questions:
RQ1: How prevalent is early test termination? In RQ1, we investigate how prevalent early test termination is in failed tests and the amount of test code that is skipped due to early test termination. In addition, we also investigate the prevalence of tests with multiple assertions, since such tests provide the soil for early test termination. Note that, if a test only has one assertion, it usually could not be the case of early test termination, since the assertion usually is the last line of the code (we have not observed any cases in our dataset where the last line is not an assertion).
RQ2: How does early test termination affect test coverage? In RQ2, we aim to investigate the impact of early test termination on the test coverage of subject under test. More specifically, we compare the test coverage in two settings: Original setting where we do not apply any approach to eliminate the early test termination; and no_stop setting, where we enforce Testearly to continue executing the rest of the code after the failed assertion(s). We introduce how we enforce Testearly to continue executing in Section III-C. By understanding this, we could provide insights to practitioners into the impact of early test termination on test coverage.
RQ3: How does early test termination affect the performance of spectrum-based fault localization? In RQ3, we aim to investigate the impact of early test termination on the performance of techniques that rely on test coverage information by comparing two settings Original and no_stop similar to RQ2. In this study, we select to study SBFL as our studied case. Note that our methodology could also be used for other techniques that rely on test coverage. By understanding this, we could provide insights to practitioners into the impact of early test termination on SBFL.
III-B Data Preparation
All | Early test termination | ||||||||
Project | Version | LOC | Avg test | Avg Tmulti | Version | Ttotal | Tearly | TearlyAssert | Cnoexecuted |
fastjson | 12 | 6K | 4,874.0 | 1691.6 (34.7%) | 4 | 82 | 32 | 8 | 8 (46.0%) |
jackson-core | 9 | 18K | 396.6 | 160.9 (40.6%) | 2 | 31 | 29 | 2 | 4.5 (15.3%) |
jfreechart | 3 | 92K | 2,178.3 | 1,472.3 (67.6%) | 2 | 4 | 3 | 3 | 19 (60.5%) |
math | 104 | 199.2K | 2,926.6 | 1,075.9 (36.8%) | 29 | 170 | 81 | 37 | 20.13 (44.7%) |
joda-time | 26 | 26K | 3,926.3 | 2,673.3 (68.1%) | 5 | 74 | 32 | 9 | 18.2 (47.1%) |
lang | 53 | 18.68K | 1,927.0 | 1,413.1 (73.3%) | 18 | 110 | 60 | 31 | 13.72 (49.6%) |
total | 207 | N/A | N/A | N/A | 60 | 471 | 237 | 90 | N/A |
For All, we present the number of versions, the average number of LOC, and the average number of tests in each version, the average number and ratio of tests with multiple assertions (Tmulti) for each project. For Early test termination, we present the number of versions, total failed tests (Ttotal), early terminated tests including the ones earlier terminated due to other reasons (Tearly), early terminated tests due to assertion failure (TearlyAssert), and the average amount of code (lines of code and ratio) in a test that is not executed due to assertion failure (Cnoexecuted) for each project.
For our experiment, we select six open-source projects from benchmark datasets Defects4J [21] and T-Evos [1]. Table I gives an overview of the descriptive statistics of the projects. Defects4J contains real bugs extracted from large open-source projects, which enables controlled experiments in software debugging and testing research [21]. Defects4J has been widely used for evaluating fault localization techniques in the research community [25, 26, 27, 28]. We selected five projects (i.e., jackson, jfreechart, commons-math (math for short), joda-time, and commons-lang (lang for short)) from Defects4J, because they are from different domains and were used as the benchmark to evaluate various SBFL in prior studies [29, 4]. To increase the diversity of our study, we further expanded our dataset with T-Evos, following an existing study [4]. T-Evos provides continuous test execution and failure data to help understand testing practices and evaluate automated testing techniques. We added one more project fastjson from T-Evos to increase the dataset diversity. In T-Evos, fastjson is the one with the highest number of test cases per commit and with the highest lines of test codes. This allows us to study early test termination on projects that have rigorous and mature testing. As some of the projects from T-Evos, such as jackson-core and jfreechart, overlap with Defects4J, we studied these projects using the versions provided by Defects4J. For each subject project, we execute the program on the selected versions provided by the benchmarks. We also verify that the number of tests is sufficient, the program builds successfully, and there is at least one failed test. In total, we collected 207 versions (i.e., commits with failed tests) from six projects for our experiment. Each project has an average number of tests ranging from 396.6 to 4,874. We summarize the basic statistics of our studied projects in Table I (All section).
III-C Methodology
Figure 1 presents the overall framework for answering our RQs. First, we identify early terminated tests due to assertion failure (i.e., Testearly) In RQ1, we perform analysis on those identified Testearly. In RQ2 and RQ3, we only focus on the versions of the projects that contain Testearly. In RQ2, we first run the test cases to collect testing information for the Original setting. Then, we apply Trycatch and Slicing to eliminate the early test termination and enforce the continuous execution of Testearly. Next, we rerun the test, in which the Trycatch and Slicing have been applied. We compare the code coverage and the performance of SBFL of Original, Trycatch, and Slicing to answer RQ2 and RQ3.
III-C1 Identifying early terminated tests due to assertion failure
The first step of our study is to identify Testearly and their corresponding versions. For this purpose, we need to build and run all the tests on every version of our studied projects. Then, we extract and analyze the failed tests for each version. Among the failed tests, we identify the ones that are Testearly by examining the reports generated by JUnit. More specifically, we analyze the testing report to examine: 1) if the terminated line is in the middle of the corresponding test; and 2) if the test failure is due to assertion failure based on our definition (Section II). To this end, we collect the log information printed during testing and extract the location of each failure, meanwhile examining whether the failure is due to assertion failure by matching keywords (e.g., “AssertionFailedError”). We identify the location of each failed assertion by analyzing the line number printed in the stack trace. Eventually, we identified 90 Testearly distributed in 60 versions (see Table I). Our analysis of the RQs is based on those 60 versions that contain at least one Testearly.
III-C2 Enforcing the execution for early terminated test (no_stop setting)
To understand the impact of early test termination, we need to compare two settings, where we eliminate the early test termination (no_stop) and without eliminating early test termination (Original). Now we introduce two approaches for no_stop and elaborate on them below.
IV Results of the research questions
RQ1: How prevalent is early test termination?
A significant portion of test cases have multiple assertions in a single test, ranging from 34.7% to 73.3% in our different studied projects. From Table I, we observe that fastjson has the lowest percentage of tests containing multiple assertions in a single test, at 34.7% on average. In addition, in the two largest projects math and lang, 36.8% and 73.3% of the tests have multiple assertions. Our findings are compatible with prior studies [44, 19], which show that a test having multiple assertions is prevalent in real-world projects and sets the stage for early test termination to occur.
Among the failed tests, a non-negligible portion (19.1%) is due to assertion failures, occurring in 29% of the studied versions. Table I shows the results of early test termination in the studied projects. Our analysis, conducted across 207 studied versions, reveals that 29% of the versions contain at least Testearly, indicating that early test termination is prevalent in the studied projects. Among the 471 failed tests observed, a significant portion (i.e., 50.3%) was terminated early, with a non-negligible portion (i.e., 19.1%) of these early terminated tests occurring due to assertion failure. To better understand the reasons for early termination that are not due to assertion failure, we manually investigated 20 samples, all of which failed due to errors in object construction or method invocation. For instance, in jfreechart, an error happens frequently when the method of is invoked, causing the test to terminate early. Future studies may investigate the effect of such failures on testing practices.
On average, 15.3% to 60.5% of the test code is not executed in Testearly due to early test termination. We examine the amount of test code that is not executed due to assertion failure. From Table I, we see that on average, 15.3% to 60.5% of code in Testearly is not executed due to assertion failure in our studied projects. The results indicate that around half of the test code in Testearly is skipped due to early test termination. For instance, in Figure LABEL:lst:example1, there are in total 10 lines of statements, and the assertion fails at line 13, leaving 2 lines of code unexecuted, which accounts for 20% of the test code.
RQ2: How does early test termination affect test coverage?
In general, early test termination negatively impacts the coverage in the production code in terms of all granularities (i.e., instruction, line, branch, and method). Table II presents the production code coverage by only Testearly of Original, Trycatch, and Slicing. We observe significant coverage improvement after eliminating early test termination. On average, applying Trycatch and Slicing improves the production code coverage by 4.5%, 5.3%, 7.8%, and 5.9% over Original in terms of instruction, line, branch, and method, respectively. For instance, in jfreechart, after eliminating early test termination, 16 (15.1%) more lines, 4.5 (25.5%) more branches, and 5 (17.5%) more methods are covered compared to Original, in which 106 lines, 17 branches, and 28.5 methods are covered. This is reasonable, since in jfreechart, an average of 60.5% of the test code is skipped due to early test termination. Intuitively, if there are fewer tests in the test suite and assertions fail earlier in the testing process, the impact of early test termination is likely to be even more pronounced.
Original | ||||
Project | Instruction | Line | Branch | Method |
math | 5,521 | 333.97 | 127.28 | 47.59 |
joda-time | 10,117.20 | 2,450.80 | 767.40 | 568.80 |
jfreechart | 388.50 | 106 | 17 | 28.50 |
jackson | 3,961 | 854 | 276.50 | 140 |
fastjson | 10,614 | 1,815 | 604.75 | 220.25 |
lang | 467.78 | 65.28 | 24.44 | 12.06 |
Trycatch and Slicing (improvement) | ||||
Project | Instruction | Line | Branch | Method |
math | 6,007.28 (486.28) | 376.24 (42.28) | 140.41 (13.14) | 55.45 (7.86) |
joda-time | 10,139.60 (22.40) | 2,455.20 (4.40) | 771.40 (3.80) | 568.80 (0) |
jfreechart | 449.50 (61) | 122 (16) | 21.50 (4.50) | 33.50 (5) |
jackson | 3,970 (9) | 856.50 (2.50) | 279.50 (3) | 140 (0) |
fastjson | 10,705.25(91.25) | 1,827 (12) | 616 (11.25) | 221.50 (1.25) |
lang | 474.33 (6.55) | 67 (1.72) | 26.06 (1.62) | 12.17 (0.11) |
Avg Imp | 4.5% | 5.3% | 7.8% | 5.9% |
* Note that the results of Slicing and Trycatch on coverage are the same. The value in “()” indicates the extra coverage achieved by Trycatch and Slicing over Original. Avg Imp stands for average improvement over different projects.
Trycatch and Slicing improve the line/branch coverage in 55% (30/60) of the studied versions compared to the Original. To provide a better view of how many versions get improved, we compare the coverage of Original, Trycatch, and Slicing for every version. Table III presents the number of versions in which the coverage gets improved, deteriorated, and tied after applying Trycatch and Slicing compared with Original in terms of four granularities, i.e., instruction, line, branch, and method. First of all, the coverage of no version gets deteriorated in all granularities, which is expected, since Trycatch and Slicing enforces Testearly to keep executing even if an assertion failure happens. In more than half of the versions, the instruction (34/60), line (33/60), and branch (33/60) coverage gets improved. In 8 out of 60 versions, the method coverage also gets improved.
In summary, our findings indicate that different assertions in a test verify different parts of the program, early test termination would miss the verification of certain production code.
Project | #Improved | #Tied | ||||||
Ins | Ln | Br | Md | Ins | Ln | Br | Md | |
math | 17 | 18 | 16 | 6 | 12 | 11 | 13 | 23 |
joda-time | 2 | 2 | 4 | 1 | 3 | 3 | 1 | 4 |
jfreechart | 1 | 1 | 0 | 1 | 1 | 1 | 2 | 1 |
jackson | 2 | 2 | 1 | 0 | 0 | 0 | 1 | 2 |
fastjson | 4 | 4 | 4 | 0 | 0 | 0 | 0 | 4 |
lang | 8 | 6 | 8 | 0 | 10 | 12 | 10 | 18 |
Total | 34 | 33 | 33 | 8 | 26 | 27 | 27 | 52 |
* Note that the number of versions that get deteriorated is zero across all studied projects. To save space, we do not show the results of deteriorated cases.
RQ3: How does early test termination affect the performance of spectrum-based fault localization?
Original | Trycatch | Slicing | ||||||||||
Project | MFR | EXAM | Top@k | MFR | EXAM | Top@k | MFR | EXAM | Top@k | |||
5 | 10 | 5 | 10 | 5 | 10 | |||||||
math | 15.2 | 0.00191 | 25.9% | 30.1% | 11.7 | 0.00188 | 30.9% | 34.3% | 9.1 | 0.00183 | 26.0% | 33.0% |
joda-time | 596.2 | 0.03306 | 0.0% | 0.0% | 562.6 | 0.03318 | 0.0% | 0.0% | 431.4 | 0.02733 | 0.0% | 0.0% |
jfreechart | 125.5 | 0.00136 | 0.0% | 0.0% | 125.5 | 0.00136 | 0.0% | 0.0% | 125.5 | 0.00136 | 0.0% | 0.0% |
jackson-core | 633.0 | 0.0383 | 0.0% | 0.0% | 631.5 | 0.03821 | 0.0% | 0.0% | 627.5 | 0.03796 | 0.0% | 0.0% |
fastjson | 7,198.3 | 0.16767 | 0.0% | 0.0% | 7,182.5 | 0.167 | 0.0% | 0.0% | 7,175.3 | 0.167 | 0.0% | 0.0% |
lang | 30.7 | 0.00237 | 6.9% | 10.23% | 27.6 | 0.00236 | 6.9% | 10.2% | 23.9 | 0.00228 | 10.7% | 14.90% |
Original | Trycatch | Slicing | ||||||||||
Project | MFR | EXAM | Top@k | MFR | EXAM | Top@k | MFR | EXAM | Top@k | |||
5 | 10 | 5 | 10 | 5 | 10 | |||||||
math | 16.4 | 0.00196 | 25.9% | 29.6% | 13.1 | 0.00190 | 32.6% | 36.1% | 10.6 | 0.00185 | 32.4% | 36.1% |
joda-time | 586.8 | 0.03051 | 0.0% | 0.0% | 583.2 | 0.03002 | 0.0% | 0.0% | 469.6 | 0.02715 | 0.0% | 0.0% |
jfreechart | 125.5 | 0.00135 | 0.0% | 0.0% | 125.5 | 0.00135 | 0.0% | 0.0% | 125.5 | 0.00135 | 0.0% | 0.0% |
jackson-core | 633.0 | 0.0383 | 0.0% | 0.0% | 627.5 | 0.03796 | 0.0% | 0.0% | 627.5 | 0.03796 | 0.0% | 0.0% |
fastjson | 7,271.0 | 0.16937 | 0.0% | 0.0% | 7232.3 | 0.16846 | 0.0% | 0.0% | 7,221.5 | 0.16821 | 0.0% | 0.0% |
lang | 25.5 | 0.00226 | 10.7% | 25.3% | 25.3 | 0.00225 | 9.3% | 24.0% | 23.7 | 0.00213 | 10.90% | 29.2% |
In general, early test termination negatively impacts the effectiveness of SBFL. Eliminating early test termination improves the performance of Ochiai and Tarantula in all evaluation metrics. For instance, Both Slicing and Trycatch can improve the performance of SBFL by 15.1% and 6.7% on average in terms of MFR, respectively, compared with Original for Ochiai. Tables IV and V present the performance of SBFL in Original, Trycatch, and Slicing in terms of MFR, EXAM, and Top@k for Ochiai and Tarantula, respectively. In general, we observe that both Slicing and Trycatch outperform Original in terms of MFR and EXAM in all projects for both Ochiai and Tarantula, except jfreechart. In other words, early test termination negatively impacts the effectiveness of SBFL. Typically, on average, Slicing improves the performance of Ochiai and Tarantula by 15.1% and 10.7% compared with Original in terms of MFR across all projects, respectively. Particularly, in the two largest projects math and lang, MFR is reduced by early test termination for both SBFLs (Ochiai and Tarantula) by a large margin. In math, MFR is deteriorated from 9.7 (Slicing) to 15.2 by 40.1% for Ochiai and is reduced from 10.6 to 16.4 by 35.4% for Tarantula, respectively. In terms of Top@k, Slicing and Trycatch improve Original in both math and lang. We do not see improvement in other projects in terms of Top@k. Although MFR has improved, none of the fault statements could be identified in the top 10 results. Similar to RQ2, as shown in Table LABEL:tab:FL_versionwise, we also observe that in a remarkable portion of the versions, Trycatch and Slicing improve the effectiveness of SBFL compared with Original, for both Ochiai and Tarantula, respectively. For instance, Slicing outperforms Original in 18 and 22 out of 60 versions for Ochiai and Tarantula, respectively.
Let’s consider an example that illustrates why eliminating early test termination (i.e., Trycatch) could potentially improve the effectiveness of SBFL in some cases. In version 72 of math, the SBFL gets improved after applying Trycatch. The test shown in Figure LABEL:lst:example1 aims to test the function shown in Figure LABEL:lst:successfulcase, which has two buggy lines located in two branches. In Original, the test terminates at the third assertion (line 13) in the test and misses the opportunity to reveal the bug (line 16) in the second branch (lines 15 - 18) in function , which is verified by the last assertion in the test (line 17) due to early test termination. While in Trycatch, all assertions are executed and both of the buggy lines are identified by SBFL.
V Discussion
V-A Overhead of our approaches
The overhead of Slicing and TryCatch is low as shown in Table VI (on average, two seconds for each system). We only applied test transformation on failed test cases with early termination and rerun them, which usually takes only milliseconds. Therefore, the overhead mainly comes from test execution. Slicing introduces extra overhead since it splits a test case with n assertions into n sub-test cases. TryCatch adds try-catch blocks to force the execution of the entire test case. In theory, TryCatch brings less overhead than Slicing, since Slicing increases the size of test cases, while TryCatch does not.
Project | Original | TryCatch | Slicing |
math | 12.96 | 13.16 | 13.29 |
joda-time | 7.02 | 7.14 | 7.23 |
jfree | 5.46 | 5.47 | 5.49 |
jackson | 4.36 | 4.41 | 4.44 |
fastjson | 43.83 | 56.62 | 56.94 |
lang | 12.50 | 12.56 | 12.67 |
V-B Implications of our findings
We recommend that developers limit each test to a single assertion. In cases where this is not feasible, use Slicing to eliminate early test termination. In the recent versions of JUnit, certain mechanisms have been developed, such as ErrorCollector starting from JUnit 4 and assertAll starting from JUnit 5, to enable the test to keep executing even if an assertion fails in the middle. Similar to Trycatch, although such mechanisms (ErrorCollector and assertAll) can avoid early test termination, our findings show that Slicing provides more accurate testing information than other mechanisms and outperforms Trycatch in locating faults. This is because Slicing disassembles a test with multiple assertions into a set of independent sub-tests, that enable the execution of every assertion and its associated code independently and provide fine-grained granularity. Therefore, we strongly encourage developers to limit each test to a single assertion whenever possible and to apply Slicing wherever multiple assertions are necessary. Although, one limitation of Slicing is that it splits one test into multiple tests and code before assertions are executed multiple times which introduces additional overhead.
We encourage future research to investigate how the industry solutions (e.g., assertAll and ErrorCollector) are applied for handling multiple assertions in a test case. Although having multiple assertions in a single test can be considered a test smell and not recommended, this is a common practice in the studied projects. It is also interesting to understand why developers do (not) use the assertAll (Junit 5) and ErrorCollector (Junit 4) mechanisms and facilitate the use of such mechanisms.
V-C Threats to Validity
Internal Validity In RQ3, we evaluate SBFL using three commonly used metrics (i.e., EXAM, top@k, and MFR). Although there might be other metrics that could be used for our task, these metrics are commonly used in evaluating SBFL techniques [2, 25]. A threat to our study is that the program transformation possibly change the behavior of test cases. To mitigate this threat, we measure the quality of transformed tests, we applied Trycatch and Slicing on passed test cases with multiple assertions and measured their code coverage. Our results show that Original, Trycatch, and Slicing have almost the same code coverage (at most 0.4 lines difference across all tests), which means the behaviour of the tests remains mostly the same. In addition, our transformation does not change the results from passed to failure. See more detailed results in our replication package [22].
External Validity Threats to external validity relate to the generalizability of our findings. In RQ3, we selected two popular SBFL techniques Ochiai and Tarantula as our subject techniques. Early test termination may have a different influence on the effectiveness of other SBFL techniques. However, our proposed framework could be applied to any techniques that rely on test coverage information. Another threat relates to our studied projects. Our findings might not be generalized to other projects. To mitigate the threat, we selected 6 projects that are in various domains from popular benchmarks Defect4J and T-Evos. Among all the 471 versions, 60 versions contain 90 Testearly. Future research may further expand the study on more datasets. In this study, we focus on early test termination due to assertion failure (only 19% of the studied early test termination), there are other reasons (e.g., error in method invitation and object construction) that could lead to early test termination. We encourage future research to investigate other reasons.
VI Related work
VI-A Understanding and improving the usage of assertion in software testing.
Xuan et al. proposed an approach to separate existing test cases into small fractions (called purified test cases) for each assertion, which is similar to Slicing, and to enhance the test oracles to further localize faults [24]. This work is the most closely related to our study. However, different from Xuan et al’s study which focuses on improving the effectiveness of fault localization by purifying test cases, our study focuses on investigating the impact of early test termination. More specifically, we investigated the prevalence of early test termination and provided empirical evidence for it. Moreover, we investigated the impact of early test termination on code coverage and fault localization by comparing it with the baseline in which we eliminated early test termination using Trycatch.
A few studies have been conducted to understand assert practice in testing [45, 44, 46, 19]. Bavota et al. studied 18 systems to analyze the distribution of test smells and show that Assertion Roulette (i.e., a code smell occurs when a test method has multiple non-documented assertions) is presented in 62% of the total JUnit classes [19]. Zamprogno et al. examined more than 30k assertions from 105 projects and also found 17.4% of the test cases have three or more assertions, which is compatible with our finding in RQ1. On the other hand, various solutions were proposed to improve the assert practice in unit testing [47, 48, 49, 50, 51]. For instance, Fraser and Arcuri proposed EvoSuite, which uses mutation testing to produce a set of assertions that maximizes the number of seeded defects in a class that can be revealed by the test cases. Watson et al. employed a Neural Machine Translation (NMT) based approach called Atlas to automatically generate meaningful assert statements providing a test method and a focal method to test the correctness of the focal method [47]. Our study differs from previous ones in that, we concentrate on examining a specific occurrence of test failure, which arises when an assert statement fails in the middle of a test case, thereby preventing the execution of code after that point.
VI-B Investigating the factors that influence testing coverage and the corresponding techniques
Masri et al. investigated the prevalence of factors that impair the effectiveness of coverage-based fault localization [52]. They found that 72% of their studied buggy versions exhibit the condition for failure was met but the program did not fail. More specifically, three conditions must be satisfied for program failure to occur: (1) the defect’s location must be reached, (2) the program’s state must become infected and (3) the infection must propagate to the output. Several studies have been done on this and demonstrate the prevalence of Coincidental correctness and its negative impact on fault detection and localization techniques [53, 54, 55, 52]. For instance, Assi et al. investigated the impact of coincidental correctness on test suite reduction (TSR), test case prioritization (TCP), and SBFL [55]. They found that TSR and TCP, the negative impact of CC was very significant.
Different from prior studies, we focus on investigating the impact of early test termination on the coverage and corresponding techniques that rely on it (i.e., SBFL). Our findings show that early test termination does negatively impact code coverage and the effectiveness of SBFL.
VII Conclusion
In this paper, we first study the prevalence of early test termination in failed tests and conduct experiments on six widely used open-source projects in software debugging and testing research from Defects4J and T-Evos benchmarks. Our findings show that 50.3% of failed tests are early terminated, with 19.1% of them caused by assertion failure. We also demonstrate that early test termination negatively impacts the code coverage across all granularities and can affect the performance of SBFL. To address this issue, we propose to use two approaches, Trycatch and Slicing, that catch assertion failures and enforce test execution to mitigate the impact of early test termination. Our evaluation indicates that Slicing and Trycatch can improve the performance of SBFL by 15.1%, 6.7% for Ochiai, and 10.66%, 3.81% for Tarantula on average in terms of MFR, respectively. Our study emphasizes the impact of early test termination on code coverage and the performance of fault localization techniques that rely on it. We also suggest that practitioners limit each test to a single assertion or use Slicing when multiple assertions are necessary.
VIII Data Availability
We made our dataset publicly available [22] to encourage future research on studying the impact of test quality on fault localization and related approaches.